(8a) Solve the ODE y - 3y = 4y (86) Solve the ODE y - 3y = 4y + 3 (9a) Solve the ODE = - 4y (9b) Solve the ODE y = -4y
Answer #1

8(a) that is y-3y = Ay y< 3y-Ay = Auxiliary equation. will be m²-3m-4 = m² Amtmat=0 mim- A) + 1 (m-A)=0 (m+1)(m-1)=0 ma -I,ora) 994 ay : Auxilary equation of be m² + 4 4. w C:f. = Opercuma C, Cosax4G sinan 9(b) PI. 4 (-8x) (D4+4) (su = (1+3) CBM) #